ROTATIONAL SPECTRA OF CYCLOPENTADIENYLTUNGSTEN TRICARBONYL HYDRIDE COMPLEX USING PULSE BEAM FOURIER TRANSFORM MICROWAVE SPECTROMETER^{a}

\underline{CHAKREE TANJAROON}, CHANDANA KARUNATILAKA, KRISTEN KECK, AND STEPHEN KUKOLICH, Department of Chemistry, University of Arizona, Tucson, AZ 85721.

Transition metal carbonyl hydride complexes based on tungsten or molybdenum have recently been explored as possible agents for hydrogenation catalysts. This is the first measurement of rotational spectra for a tungsten hydride complex. The asymmetric-top spectra for the four tungsten isotopomers of cyclopentadienyltungsten tricarbonyl hydride complex were measured in the frequency range of 5-12 GHz using a Flygare-Balle type of spectrometer. The spectra have been assigned and rotational parameters have been derived from the least-squares fits. The preliminary fit results suggest that the complex behaves nearly like a rigid rotor. The work is in progress to obtain spectra for more isotopic species to solve the gas phase structure and obtain the hydrogen bond length of this complex.

aSupported by THE NATIONAL SCIENCE FOUNDATION